# A Symmetric Multigrid-Preconditioned Krylov Subspace Solver for Stokes Equation

### v1

@@ Abstract
Numerical solution of discrete PDEs corresponding to saddle point problems is highly relevant to physical systems such as Stokes flow. However, scaling up numerical solvers for such systems is often met with challenges in efficiency and convergence. Multigrid is an approach with excellent applicability to elliptic problems such as the Stokes equations, and can be a solution to such challenges of scalability and efficiency. The degree of success of such methods, however, is highly contingent on the design of key components of a multigrid scheme, including the hierarchy of discretizations, and the relaxation scheme used. Additionally, in many practical cases, it may be more effective to use a multigrid scheme as a preconditioner to an iterative Krylov subspace solver, as opposed to striving for maximum efficacy of the relaxation scheme in all foreseeable settings. In this paper, we propose an efficient symmetric multigrid preconditioner for the Stokes Equations on a staggered finite-difference discretization. Our contribution is focused on crafting a preconditioner that (a) is symmetric indefinite, matching the property of the Stokes system itself, (b) is appropriate for preconditioning the SQMR iterative scheme freund1994new , and (c) has the requisite symmetry properties to be used in this context. In addition, our design is efficient in terms of computational cost and facilitates scaling to large domains.
@@
Stokes equations, Geometric multigrid, Preconditioner, Krylov subspace solver, Staggered finite-difference method
PACS:
$0$, $1$
MSC:
$0$, $1$
@@
{graphicalabstract}
[Uncaptioned image][Uncaptioned image]
{highlights}
We design effective symmetric smoother for both the distributive smoother and Vanka smoother, combining them to achieve a balance between convergence rates and computational time per iteration.
@@ 1 Introduction
Saddle-point problems arise in many fields such as fluid dynamics elman2002preconditioners ; de2007two , structure mechanics farhat1992saddle ; franceschini2019block and magnetohydrodynamics phillips2016block . As modern computational devices develop, a requirement for solving large-scale problems arises. Multigrid methods trottenberg2000multigrid are well-known for their potential scalability in handling large-scale problems with the advantage of linear time and space complexity. Although the convergence rates of multigrid methods are independent of problem sizes, the design and optimization of multigrid components, such as relaxation schemes, can significantly impact convergence rates. The relaxation scheme, often referred to as a smoother in multigrid, comprises various classical techniques, including the distributive smoother oosterlee2006multigrid , Uzawa smoother elman1994inexact , Braess-Sarazin smoother braess1997efficient and Vanka smoother vanka1986block . The distributive smoother transforms the original system equations into a right-preconditioned system, aiming to improve properties like condition numbers. It has been studied for solving problems like the Stokes equations oosterlee2006multigrid ; bacuta2011new ; wang2013multigrid and the Oseen equations chen2015multigrid . The Uzawa smoother transforms indefinite systems into positive definite ones by using Schur complement, with applications in solving the Stokes equations maitre1985fast and poroelasticity equations luo2017uzawa . The Braess-Sarazin smoother is a variant of the pressure correction steps in SIMPLE-type algorithms patankar1983calculation . It also relies on the approximation of Schur complement and has been introduced for tackling challenges such as the Stokes equations braess1997efficient ; he2018local and magnetohydrodynamic equations adler2016monolithic . In contrast to the previous three smoothers, the Vanka smoother focuses on solving local overlapping saddle-point problems instead of the entire saddle-point system. While the Vanka smoother shows high efficiency in each iteration, it comes with higher per-iteration costs compared to other smoothing methods. It has proven effective in the Stokes equations vanka1986block ; saberi2022restricted and poroelasticity equations franco2018multigrid . In addition to the choice of smoother, the convergence of multigrid also depends on factors like grid-operators niestegge1990analysis , discretization methods brandt2011multigrid , and even on some problem-specific parameters parsons1990multigrid .
@@
Considering the sensitivity of multigrid methods, they can be more effective when used as a preconditioner to an iterative Krylovs subspace solver. The multigrid-preconditioned conjugate gradient (MGPCG) method tatebe1993multigrid is the prototypical approach for solving positive definite systems such as Poisson’s equation. However, for saddle-point problems, a more general Krylov subspace solver is required. The generalized minimal residual (GMRes) method saad1986gmres emerges as the most common choice in this context, promising convergence for any asymmetric indefinite system. The multigrid-preconditioned GMRes has been employed in addressing transport equations oliveira1998preconditioned , advection-diffusion equations oosterlee1996use ; ramage1999multigrid and Navier-Stokes equations anselmann2023efficiency . Despite of the indefiniteness of saddle-point problems, symmetry prevails in many situations such as the Stokes equation, Oseen equations and Helmholtz equation. The minimal residual (Minres) paige1975solution method, designed for symmetric indefinite systems, consumes both less memory and less computational time than GMRes in one iteration while maintaining similar convergence rates. It appears to be a more promising option in these symmetric scenarios. However, Minres has the limitation in that it can only be paired with positive definite preconditioners, making it incompatible with multigrid preconditioning. For this reason, we investigate the symmetric quasi-minimal residual (SQMR) method freund1994new , a variant of MINRES with identical work and storage requirements. SQMR supports symmetric indefinite preconditioners, opening up opportunities for the use of multigrid as the preconditioner. To the best of our knowledge, no multigrid-preconditioned SQMR has been developed for symmetric saddle-point problems.
@@
Since SQMR requires symmetric indefinite preconditioners, the smoothing step of multigrid must also be symmetric. Inspired by the symmetric Gauss-Seidel method, we design an effective symmetric distributive smoother and Vanka smoother in this work. The main idea behind the symmetric smoother involves relaxation in some specific order, followed by relaxation in the exact reverse order. Our primary focus is on the multigrid-preconditioned SQMR method, utilizing a staggered finite-difference discretization of the Stokes equations as the model problem. The main contributions of this work can be summarized as follows:
@@
    We design effective symmetric smoother for both the distributive smoother and Vanka smoother, combining them to achieve a balance between convergence rates and computational time per iteration.
@@
The remaining structure of the paper is as follows: In section 2, we introduce the finite-difference discretization of the Stokes equation and explain how we apply multigrid. In section 3, we discuss how we ensure the symmetry of the distributive smoother and Vanka smoother. In section 4, we propose our approach to use multigrid as a preconditioner for SQMR. In section 5, we describe our discrete domain design used for boundary conditions. In section 6, we compare the performance of our multigrid-preconditioned solver with the classical multigrid method on both 2D and 3D benchmarks. Finally, we draw our conclusions in section 7.
2 Multigrid for the Stokes equation
In this work, we consider the following Stokes equations, which apply to both 2D and 3D scenarios:
@@ 2 Multigrid for the Stokes equation
where $\mathbf{u}=(u,v)^{T}$ (2D) or $\mathbf{u}=(u,v,w)^{T}$ (3D) is the fluid velocity vector field, $p$ is the pressure scalar field, $\mathbf{f}$ is the body force vector field and $\eta$ is the fluid viscosity, together with a suitable boundary condition.
@@
Then we discuss the skeleton of our multigrid method. When it comes to balancing computational time and convergence rates, we have various multigrid methods to choose from, such as V-cycle, W-cycle and F-cycle stuben2006multigrid . For efficiency, we opt for V-cycle, which runs faster per iteration but converges slightly slower than the others. Since multigrid is intended to be used as a preconditioner, V-cycle is an adequate design. The recursive V-cycle procedure is outlined in Alg. 1, where the superscript $h$ denotes discretization at grid size $h$. It is worth mentioning that we rediscretize $\mathbf{L}^{2h}$ at a coarser level rather than using Galerkin coarsening brandt2011multigrid , using finite-difference discretization. The locations of $u$ and $p$ components of 2D fine grid and coarse grid are shown in Fig. 2. It can be analogized to $v$ component and 3D cases.
Figure 2: 2D discretization on fine and coarse level
As we mentioned above, the restriction operator $\mathbf{R}$ and prolongation operator $\mathbf{P}$ can also influence convergence rates for the Stokes equations. Some common combinations have been investigated by Niestegge niestegge1990analysis . To maintain the symmetry (discussed further later), we must use the transposed operators for $\mathbf{R}$ and $\mathbf{P}$, meaning $\mathbf{P}=c\mathbf{R}^{T}$ with $c$ is typically 4 for 2D and 8 for 3D. To balance the symmetry and the convergence rates, we use the bilinear interpolation for velocities and constant per-cell interpolation for pressures as the prolongation operator. The restriction operator is exactly transposed of the prolongation operator divided by $c$. Fig. 3 illustrates how we prolongate on $u$ component and $p$ component in 2D cases.
Figure 3: 2D Prolongation Operator
@@
7:         $\mathbf{x}^{h}\leftarrow\text{Smoothing}(\mathbf{L}^{h},\mathbf{x}^{h}_{0},\mathbf{b}^{h})$
@@
14:         $\mathbf{x}^{h}\leftarrow\text{Smoothing}(\mathbf{L}^{h},\mathbf{x}^{h},\mathbf{b}^{h})$
@@
Algorithm 1 Recursive V-cycle
3 Smoothing step and symmetry
As discussed in MGPCG tatebe1993multigrid , the symmetry of the V-cycle can be reduced to the symmetry of the smoothing step. A smoothing step can be formulated as
@@ 3 Smoothing step and symmetry
where $k$ denotes the iteration step and $\mathbf{S}$ is the smoother operator. For example, classical Gauss-Seidel smoothing can be expressed as $\mathbf{S}_{GS}=\mathbf{G}^{-1}$ with $\mathbf{G}$ being the lower triangular component of system matrix $\mathbf{L}$. Similarly, the weighted Jacobi method can be written as $\mathbf{S}_{WJ}=\omega\mathbf{D}$, where $\omega$ is the damping factor and $\mathbf{D}$ is the diagonal component of $\mathbf{L}$. We can observe that $\mathbf{x}^{k}$ must be a linear transformation of $\mathbf{b}$ when $\mathbf{x}^{0}=\mathbf{0}$, implying $\mathbf{x}^{k}=\mathbf{C}\mathbf{b}$. The smoothing step is symmetric to design a multigrid preconditioner if and only if $\mathbf{C}$ is symmetric. This simplification occurs when we employ the same smoother before and after the rediscretization step, and use the transposed operator for $\mathbf{R}$ and $\mathbf{P}$. In this section, we discuss the design of symmetric operators for both the distributive smoother and Vanka smoother. We also explore how these operators can be combined while maintaining symmetry.
@@ 3.1 Symmetric distributive smoother
The idea of the distributive smoother can be viewed as a transformation of the original system $\mathbf{L}\mathbf{x}=\mathbf{b}$ into a right-preconditioned system $\mathbf{L}\mathbf{M}\mathbf{y}=\mathbf{b}$, where $\mathbf{M}\mathbf{y}=\mathbf{x}$. For the Stokes equations, the transformed system can be formulated as
@@
$\begin{split}\mathbf{L}\mathbf{M}&=\begin{pmatrix}\mathbf{A}&\mathbf{B}^{T}\\ \mathbf{B}&\mathbf{0}\\ \end{pmatrix}\begin{pmatrix}\mathbf{I}&-\mathbf{B}^{T}\\ \mathbf{0}&\eta\mathbf{B}\mathbf{B}^{T}\\ \end{pmatrix}=\begin{pmatrix}-\eta\Delta^{h}&\nabla^{h}\\ -\nabla\cdot^{h}&\mathbf{0}\\ \end{pmatrix}\begin{pmatrix}\mathbf{I}&-\nabla^{h}\\ \mathbf{0}&-\eta\nabla\cdot^{h}\nabla^{h}\\ \end{pmatrix}\\ &=\begin{pmatrix}-\eta\Delta^{h}&\eta\Delta^{h}\nabla^{h}-\eta\nabla^{h}(\nabla\cdot^{h}\nabla^{h})\\ -\nabla\cdot^{h}&\nabla\cdot^{h}\nabla^{h}\\ \end{pmatrix}=\begin{pmatrix}-\eta\Delta^{h}&\mathbf{0}\\ -\nabla\cdot^{h}&\nabla\cdot^{h}\nabla^{h}\\ \end{pmatrix}\\ \end{split}$ (5)
The top right block is $\mathbf{0}$ due to the commutativity of $\Delta$ and $\nabla$ operators, which is a property that Stokes equation specifically holds under a finite difference discretization. It’s important to note that when alternative discretization methods like finite element are employed, the top-right block may not be exact zero but only asymptotically zero up to truncation error. Furthermore, we should also notice that the matrix can be converted into discrete operators only when the degrees of freedom are in the interior region, free from the constraints of the boundary conditions. The introduction of the distributive matrix $\mathbf{M}$ is motivated by the fact that $\mathbf{L}\mathbf{M}$ forms a block upper triangular matrix with the Laplacian operator on the diagonal block. This structure allows us to apply classical smoothing methods like Gauss-Seidel or weighted Jacobi on $\mathbf{L}\mathbf{M}$ matrix, leveraging their effectiveness on the Laplacian operator. From that, $\delta\mathbf{y}$ is obtained to compute corrections $\delta\mathbf{x}=\mathbf{M}\delta\mathbf{y}$. The smoothing operator of distributive Gauss-Seidel (DGS) and distributive weighted Jacobi (DWJ) can be represented as $\mathbf{S}_{DGS}=\mathbf{M}\mathbf{\tilde{G}}^{-1}$ and $\mathbf{S}_{DWJ}=\omega\mathbf{M}\mathbf{\tilde{D}}^{-1}$ where $\mathbf{\tilde{G}}$ and $\mathbf{\tilde{D}}$ is the upper triangular and diagonal component of system matrix $\mathbf{L}\mathbf{M}$. Instead of computing the entire $\mathbf{L}\mathbf{M}$ matrix and applying Gauss-Seidel or weighted Jacobi, this process can be simplified by storing the diagonal part of $\mathbf{L}\mathbf{M}$ and using the connection $\mathbf{x}=\mathbf{M}\mathbf{y}$. Alg. 2 provides a detailed example of the simplification of DGS smoothing. A similar approach can be employed for DWJ smoothing.
@@
Algorithm 2 Distributive Gauss-Seidel Smoothing
As we mentioned above, we aim to design a symmetric smoother. However, the symmetry of distributive smoothing is hindered by right-preconditioning. To maintain symmetry, we introduce left-preconditioning of the distributive smoother as follows
@@
It’s obvious that $\mathbf{M}^{T}\mathbf{L}=(\mathbf{LM})^{T}$ due to the symmetry of $\mathbf{L}$. By combining right DWJ with left DWJ or right DGS in a specific order with left DGS in the exact reverse order, the smoother can recover its symmetry. A proof for symmetric DGS can be found in A, and a similar proof applies to DWJ. Despite the efficiency of the distributive smoother, its convergence heavily relies on the commutativity of $\Delta$ and $\nabla$ operators in the interior region. This property doesn’t hold near the boundary, potentially leading to divergence. Therefore, it’s essential to employ more robust smoother near the boundary.

### v3

@@ Abstract
Numerical solution of discrete PDEs corresponding to saddle point problems is highly relevant to physical systems such as Stokes flow. However, scaling up numerical solvers for such systems is often met with challenges in efficiency and convergence. Multigrid is an approach with excellent applicability to elliptic problems such as the Stokes equations, and can be a solution to such challenges of scalability and efficiency. The degree of success of such methods, however, is highly contingent on the design of key components of the multigrid scheme, including the hierarchy of discretizations, and the relaxation scheme used. Additionally, in many practical cases, it may be more effective to use a multigrid scheme as a preconditioner to an iterative Krylov subspace solver, as opposed to striving for maximum efficacy of the relaxation scheme in all foreseeable settings. In this paper, we propose an efficient symmetric multigrid preconditioner for the Stokes Equations on a staggered finite difference discretization. Our contribution is focused on crafting a preconditioner that (a) is symmetric indefinite, matching the property of the Stokes system itself, (b) is appropriate for preconditioning the SQMR iterative scheme [1], and (c) has the requisite symmetry properties to be used in this context. In addition, our design is efficient in terms of computational cost and facilitates scaling to large domains.
@@
Stokes equations , Geometric multigrid , Preconditioner , Krylov subspace solver , Staggered finite difference method
@@ 1 Introduction
Saddle-point problems arise in many fields such as fluid dynamics [2, 3], structure mechanics [4, 5] and magnetohydrodynamics [6]. As modern computational platforms advance, the demand for solving large-scale problems becomes more pronounced. However, the performance potential of modern computing platforms is often hindered by existing algorithms, which may lack the necessary efficiency in design. Multigrid methods [7] are designed specifically for their potential scalability in handling large-scale problems with the advantage of linear time and space complexity in principle. Although the convergence rates of multigrid methods are – in the best-case scenario – independent of problem sizes, the design and optimization of multigrid components, such as relaxation schemes, can significantly impact this property. The relaxation scheme, often referred to as a smoother in multigrid, can be designed according to various classical techniques, including the distributive smoother [8], Uzawa smoother [9], Braess-Sarazin smoother [10] and Vanka smoother [11]. The distributive smoother transforms the original equations into a right-preconditioned system, aiming to improve properties such as conditioning or numerical structure that facilitates the applicability of the relaxation scheme. This has been studied for solving problems like the Stokes equations [8, 12, 13] and the Oseen equations [14]. The Uzawa smoother transforms indefinite systems into positive definite formulations by using Schur complement, with applications in solving the Stokes equations [15] and poroelasticity equations [16]. The Braess-Sarazin smoother is a variant of the pressure correction steps in SIMPLE-type algorithms [17]. It also relies on the approximation of Schur complement and has been introduced for tackling challenges such as the Stokes equations [10, 18] and magnetohydrodynamic equations [19]. In contrast to the previous three smoothers, the Vanka smoother focuses on solving local overlapping saddle-point problems and updating several local degrees of freedom collectively. While the Vanka smoother is highly effective in reducing local residuals, it comes with higher per-iteration costs compared to other smoothing methods. It has proven effective in the Stokes equations [11, 20] and poroelasticity equations [21]. In addition to the choice of smoother, the convergence of multigrid also depends on factors like grid-operators [22] and discretization methods [23].
@@
Considering the sensitivity of multigrid methods, they are often more effective when used as a preconditioner to an iterative Krylov subspace solver. The multigrid-preconditioned conjugate gradient (MGPCG) method [24] is the prototypical approach for solving positive definite systems such as the Poisson equation. However, for saddle-point problems, a more general Krylov subspace solver is required, due to the indefiniteness of the discretized equations. The generalized minimal residual (GMRes) method [25] emerges as the most common choice in this context, promising convergence for any asymmetric indefinite system. The multigrid-preconditioned GMRes has been employed in addressing transport equations [26], advection-diffusion equations [27, 28] and Navier-Stokes equations [29]. Despite the indefiniteness of saddle-point problems, symmetry is present as a property in many situations such as the Stokes equation, Oseen equations and Helmholtz equation. The minimal residual (Minres) [30] method, designed for symmetric indefinite systems, consumes both less memory and less computational time than GMRes per iteration while maintaining similar convergence rates. Thus, it is a more appropriate option in such symmetric scenarios. However, Minres has the limitation that it can only be paired with positive definite preconditioners, making it incompatible with multigrid preconditioning (which for a problem like ours would yield a symmetric indefinite preconditioner). For this reason, we turn our attention to the symmetric quasi-minimal residual (SQMR) method [1], a variant of MINRES with identical work and storage requirements. SQMR supports symmetric indefinite preconditioners, opening up opportunities for the use of multigrid as the preconditioning scheme. To the best of our knowledge, no multigrid-preconditioned SQMR has been developed for symmetric saddle-point problems.
@@
Since SQMR requires a symmetric (indefinite) preconditioner, care must be taken to preserve the symmetry property if multigrid is used in this context. This requires a number of design choices, many of which are trivial to implement (e.g. making sure that the prolongation and restriction are adjoint operators), but also requires a more delicate design of the relaxation scheme, so to not hinder the symmetry property. Inspired by the steps needed to preserve symmetry within the Gauss-Seidel method, we design an effective symmetric distributive smoother and Vanka smoother in our work. The main idea behind the design of a symmetric smoother involves relaxation in a specific order, followed by relaxation in the exact reverse order. For distributive relaxation, this also needs to be paired with a combination of left- and right-preconditioning in order to preserve symmetry. Our primary focus is on the multigrid-preconditioned SQMR method, utilizing a staggered finite difference discretization of the Stokes equations as the model problem. The main contributions of this work can be summarized as follows:
@@
    We design two multigrid smoothers, based on distributive relaxation and Vanka relaxation respectively, taking care to preserve the symmetry of the operators. We combine these relaxation operators selectively, to achieve a balance between convergence rates and computational time per iteration.
@@
The remaining structure of the paper is as follows: In section 2, we introduce the finite difference discretization of the Stokes equation and explain how we apply multigrid. In section 3, we discuss how we ensure the symmetry of the distributive smoother and Vanka smoother. In section 4, we propose our approach to use multigrid as a preconditioner for SQMR. In section 5, we describe our discrete domain design used for setting boundary conditions. In section 6, we compare the performance of our multigrid-preconditioned solver with the classical multigrid method on both 2D and 3D benchmarks, and also benchmark against un-preconditioned approaches. In section 7, we present additional experiments to understand how problems of different resolutions and conditioning, and multigrid cycle schemes can affect our results. Finally, we draw our conclusions in section 8.
2 Multigrid for the Stokes equations
In this work, we consider the Stokes equations as written below, which apply to both 2D and 3D scenarios:
@@ 2 Multigrid for the Stokes equations
where $\mathbf{u}=(u,v)^{T}$ (2D) or $\mathbf{u}=(u,v,w)^{T}$ (3D) is the fluid velocity vector field, $p$ is the scalar pressure field, $\mathbf{f}$ is the body force vector field and $\eta$ is the fluid viscosity, together with suitable boundary conditions. In this work, we treat the viscosity parameter $\eta$ as spatially constant.
@@
We proceed to describe the details and composition of our multigrid method. For our approach, which ultimately will be used to craft a preconditioner for an iterative Krylov solver, we adopt a V-cycle paradigm [31]; while our contributions related to symmetric relaxation schemes are equally applicable to other types of multigrid cycles such as W-cycle.
The recursive V-cycle procedure is outlined in Alg. 16, where the superscript $h$ denotes the discretization grid size at every level of the multigrid hierarchy. We highlight that the operators $\mathbf{L}^{2h},\mathbf{L}^{4h},\ldots$ at coarser multigrid levels are constructed in our work based on a re-discretization of a coarsened description of the domain at successively coarser resolutions, as opposed to an algebraic Galerkin coarsening approach [23] for example; for details of our domain coarsening approach see section 5. The locations of $u$ and $p$ components of the 2D fine coarse grids are shown in Fig. 2. The placement of variables and grids is analogous in 3D, with an 8-to-1 cell subdivision, and retaining face-centered velocities and cell-centered pressures.
Figure 2: 2D discretization on fine and coarse level; just the horizontal component $u$ of the velocity field $\mathbf{u}=(u,v)$ is illustrated on the left, the placement of vertical velocity components $v$ is analogous, but on faces perpendicular to the $y$-axis.
As we mentioned above, the restriction operator $\mathbf{R}$ and prolongation operator $\mathbf{P}$ can also influence convergence rates for the Stokes equations. Some common combinations have been investigated by Niestegge [22]. Although this is not strictly necessary if multigrid is to be used as a solver outright, if we seek to use it as a symmetric preconditioner, we must select restriction and prolongation operators $\mathbf{R}$ and $\mathbf{P}$ that are the transpose of each other up to scaling, meaning that $\mathbf{P}=c\mathbf{R}^{T}$ with $c$ typically being 4 for 2D and 8 for 3D. This scale factor (which does not impact the symmetry property) is typically selected to make both the prolongation and the restriction operators a partition of unity. To balance symmetry and the convergence rates, we use bilinear interpolation for velocities and constant per-cell interpolation for pressures as the prolongation operator. The restriction operator is thus defined as the transposed prolongation operator divided by $c$. Fig. 3 illustrates how the restriction stencil on $u$ component (the horizontal velocity component, the vertical one being analogous and symmetrical) and $p$ component in 2D cases.
Figure 3: 2D restriction operator; just the horizontal component $u$ of the velocity field $\mathbf{u}=(u,v)$ is illustrated on the left, the placement of vertical velocity components $v$ is analogous, but on faces perpendicular to the $y$-axis.
Alg. 1 Recursive V-cycle
@@
7:         $\mathbf{x}^{h}\leftarrow\text{SMOOTH}(\mathbf{L}^{h},\mathbf{x}^{h}_{0},\mathbf{b}^{h})$
@@
14:         $\mathbf{x}^{h}\leftarrow\text{SMOOTH}(\mathbf{L}^{h},\mathbf{x}^{h},\mathbf{b}^{h})$
@@
3 Design of smoothing operator and preserving symmetry
Alg. 16 details the procedure implementing a multigrid V-cycle. Iterative application of this cycle can be used to materialize an iterative solver, if used outright and not in the context of preconditioning. The same cycle can also be used to design a preconditioner appropriate for Krylov Subspace iterative methods. Tatebe [24] presented this process in the context of multigrid-preconditioned Conjugate Gradients, but the process by which the preconditioner is applied is more general across Krylov subspace solvers, and as follows: The preconditioner is intended to be a matrix $\mathbf{W}^{\dagger}\approx\mathbf{L}^{-1}$ approximating the inverse of the system matrix $\mathbf{L}$. For our purposes the preconditioner must be a symmetric matrix, although we will use it in the context of a Krylov subspace solver that allows it to be indefinite; let us highlight that in the context of Stokes equations, an excellent preconditioner would easily be expected to be symmetric indefinite, as the Stokes equations themselves are.
As detailed in section 4, a matrix-vector multiplication $\mathbf{t}\leftarrow\mathbf{W}^{\dagger}\mathbf{s}$ is implicitly implemented as $\mathbf{t}\leftarrow\textsc{V-cycle}(\mathbf{L}^{h},\mathbf{b}^{h}\leftarrow\mathbf{s},\mathbf{x}_{0}^{h}\leftarrow 0,n)$ via a call to the V-Cycle routine. Note that we have used the input vector $\mathbf{s}$ in the place of the right-hand-side vector $\mathbf{b}^{h}$, and used a zero initial guess $\mathbf{x}_{0}^{h}$ (the latter is essential in making sure the resulting operator is truly linear, as opposed to affine). Tatebe [24] identified the requirements for an operator $\mathbf{W}^{\dagger}$ constructed in this way to be symmetric. Two of the requirements are enforced by design, namely that the prolongation and restriction operators are transposed (up to scaling) of each other, and that the operator applied at the coarsest level of discretization is symmetric (here, by virtue of being an exact solution, i.e. an exact inverse of a symmetric matrix).
The more delicate requirement for the symmetry of $\mathbf{W}^{\dagger}$ is that the smoothing operators at the restriction and prolongation phases of the V-cycle (lines 7 and 14 of Alg. 16) are transposes of each other. A straightforward way to enforce that would be to ensure that these smoothing operators are symmetric in their own right. The smoothing operator can be formalized as:
@@ 3 Design of smoothing operator and preserving symmetry
where $k$ denotes the iteration step and $\mathbf{S}$ is the iteration matrix. For example, classical Gauss-Seidel smoothing can be expressed as $\mathbf{S}_{GS}=\mathbf{G}^{-1}$ with $\mathbf{G}$ being the lower triangular component of system matrix $\mathbf{L}$. Similarly, the weighted Jacobi method can be written as $\mathbf{S}_{WJ}=\omega\mathbf{D}^{-1}$, where $\omega$ is the damping factor and $\mathbf{D}$ is the diagonal component of $\mathbf{L}$.
Equation 3 can be unrolled to yield $\mathbf{x}^{k}=\mathbf{C}\mathbf{b}+\mathbf{N}\mathbf{x}_{0}$. As detailed by Tatebe [24], the symmetry of the smoothing operator for the purposes of crafting a symmetric preconditioner reduces to the symmetry of the matrix $\mathbf{C}$. We focus in our present work on the case where the same smoothing process is used in both the restriction and prolongation phases of the V-cycle; should it be an option to use different smoothing procedures at each phase, the corresponding $\mathbf{C}$ matrix for the two should be the transpose of each other; we do not focus on this case in our work though.
In the remainder of this section, we detail the steps that need to be taken such that this matrix $\mathbf{C}$ can be guaranteed to be symmetric, both when we use a distributive relaxation paradigm or a Vanka relaxation. We also explore how these operators can be combined while maintaining symmetry.
@@ 3.1 Symmetric distributive smoother
The design principle of a distributive smoother is that it can be viewed as a transformation of the original system $\mathbf{L}\mathbf{x}=\mathbf{b}$ into a right-preconditioned system $\mathbf{L}\mathbf{M}\mathbf{y}=\mathbf{b}$, where $\mathbf{M}\mathbf{y}=\mathbf{x}$. The hope is that the numerical properties of the combined matrix $\mathbf{LM}$ are more favorable for applying a standard relaxation procedure, rather than using the same procedure directly on $\mathbf{L}$. Furthermore, any relaxation on the system $\mathbf{LM}\mathbf{y}=\mathbf{b}$ can be implicitly emulated on the original system, without having to enact a change of variables from $\mathbf{x}$ to $\mathbf{y}$ beforehand. Specifically, any update $y_{i}\leftarrow y_{i}+\delta$ to one of the components of $\mathbf{y}$ can be emulated as $\mathbf{x}\leftarrow\mathbf{x}+\delta\mathbf{M}\mathbf{e}_{i}=\mathbf{x}+\delta\mathbf{m}_{i}$ on the $\mathbf{x}$, where $\mathbf{e}_{i}$ is the $i$-th basis vector (or, equally, $\mathbf{m}_{i}$ is the $i$-th column of $\mathbf{M}$)
For the Stokes equations, the transformed system can be formulated as
@@
$\begin{split}\mathbf{L}\mathbf{M}&=\begin{pmatrix}\mathbf{A}&\mathbf{B}^{T}\\ \mathbf{B}&\mathbf{0}\\ \end{pmatrix}\begin{pmatrix}\mathbf{I}&-\mathbf{B}^{T}\\ \mathbf{0}&\eta\mathbf{B}\mathbf{B}^{T}\\ \end{pmatrix}=\begin{pmatrix}-\eta\Delta^{h}&\nabla^{h}\\ -\nabla\cdot^{h}&\mathbf{0}\\ \end{pmatrix}\begin{pmatrix}\mathbf{I}&-\nabla^{h}\\ \mathbf{0}&-\eta\nabla\cdot^{h}\nabla^{h}\\ \end{pmatrix}\\ &=\begin{pmatrix}-\eta\Delta^{h}&\eta\Delta^{h}\nabla^{h}-\eta\nabla^{h}\Delta^{h}\\ -\nabla\cdot^{h}&\Delta^{h}\\ \end{pmatrix}=\begin{pmatrix}-\eta\Delta^{h}&\mathbf{0}\\ -\nabla\cdot^{h}&\Delta^{h}\\ \end{pmatrix}\\ \end{split}$ (5)
Here the superscript $h$ indicates the operator is discretized by a finite difference method with a grid size of $h$. The top right block is $\mathbf{0}$ due to two facts. First, $\nabla^{h}\cdot\nabla^{h}$ is equivalent to $\Delta^{h}$, which is a classical 5(2D) or 7(3D)-point Laplacian stencil in finite difference method. Secondly, the $\Delta$ and $\nabla$ operators are commutative, which is a property that Stokes equation specifically preserves under a finite difference discretization. It’s important to note that when alternative discretization methods like Finite Elements are employed, the top-right block may not be exactly zero but only asymptotically zero up to truncation error. Furthermore, we should also note that the discrete matrix structure illustrated in equation 5 is specific to the interior of the computational domain, since the presence of boundary equations could affect it (e.g. the triangular property of the combined matrix).
The introduction of the distribution matrix $\mathbf{M}$ is motivated by the fact that $\mathbf{L}\mathbf{M}$ forms a block upper triangular matrix with the Laplacian operator on the diagonal block. This structure allows us to apply classical smoothing methods like Gauss-Seidel or weighted Jacobi on $\mathbf{L}\mathbf{M}$ matrix, leveraging their effectiveness on the Laplacian operator. The smoother for each component block will be applied in sequence. One way to appreciate this process would be as follows: If instead of a relaxation procedure, we applied a full solution for the Laplacian at each diagonal block (or, equivalently, if we smoothed to full convergence), iterating from one block component to the next would have been a block forward-substitution process. In practice, following the same relaxation schedule is a highly effective overall smoother, even if only a small number of relaxation steps are used at each block.
From that, $\delta\mathbf{y}$ is obtained to compute corrections $\delta\mathbf{x}=\mathbf{M}\delta\mathbf{y}$. The smoothing operator of, respectively, distributive Gauss-Seidel (DGS) and distributive weighted Jacobi (DWJ) can be represented as $\mathbf{S}_{DGS}=\mathbf{M}\mathbf{\tilde{G}}^{-1}$ and $\mathbf{S}_{DWJ}=\omega\mathbf{M}\mathbf{\tilde{D}}^{-1}$ where $\mathbf{\tilde{G}}$ and $\mathbf{\tilde{D}}$ is the upper triangular and diagonal component of system matrix $\mathbf{L}\mathbf{M}$. Instead of computing the entire $\mathbf{L}\mathbf{M}$ matrix and applying Gauss-Seidel or weighted Jacobi, this process can be simplified by storing the diagonal part of $\mathbf{L}\mathbf{M}$ and using the connection $\mathbf{x}=\mathbf{M}\mathbf{y}$. Alg. 2 provides a detailed example of the simplification of DGS smoothing. A similar approach can be employed for DWJ smoothing.
@@
Alg. 2 Distributive Gauss-Seidel Smoothing
As mentioned above, we aim to design a symmetric smoother. However, the symmetry of distributive smoothing is hindered by right-preconditioning; we can easily show that applying an otherwise symmetric smoother (like Gauss-Seidel) to the non-symmetric distributed operator $\mathbf{LM}$ would compromise symmetry. In order to restore the symmetry property, we introduce left-preconditioning of the distributive smoother as follows
@@
It’s obvious that $\mathbf{M}^{T}\mathbf{L}=(\mathbf{LM})^{T}$ due to the symmetry of $\mathbf{L}$. Our combined, symmetric distributive smoother operates as follows: In a first sweep, we choose a given traversal order of the degrees of freedom, and apply DGS according to Equation 4 at each point. In a second sweep, we use exactly the reverse traversal order of the first step, and apply DGS, but in accordance with the left-preconditioned system in Equation 6. The smoother constructed in this way is symmetric; a full proof of this fact is given in A, and a similar proof applies to DWJ. Despite the efficiency of the distributive smoother, its convergence heavily relies on the commutativity of $\Delta$ and $\nabla$ operators in the interior region. This property does not discretely hold near the boundary, where this smoothing procedure is no longer applicable. Thus, we employ a different, more general and robustly convergent (albeit not as computationally efficient) smoother in the vicinity of the domain boundary.